# Widespread disruption of resonant chains during protoplanetary disk dispersal

### v1

@@ 4.1 Dependance on model parameters
Figure 5 shows the evolution of a pair of $6M_{\oplus}$ planets in three different scenarios. In each case, the accretion disk evolution is the same, but we have scaled the form of $a_{R}$, given in equation (4), by a constant factor. As discussed above, this mimics variations in either the stellar magnetic field strength, or the degree of field compression by the inflow ($n$). Thus the black curve in Figure 5 shows the default case, while the blue curve shows the case where $a_{T}$ is a factor 0.7 smaller (corresponding to a weaker magnetospheric influence) and the red case represents the case where $a_{T}$ is a factor 1.3 larger (corresponding to a stronger magnetospheric influence).
@@
The final separations also depend on the model for the evolution of the accretion rate and surface density of the disk. Our default model assumes an evolution characterised by two exponentials, chosen to mimic the slow decay of a viscously evolving disk that accelerates once the inner disk has been decoupled from the outer disk by photoevaporation. Figure 6 shows how the evolution of a pair of $6M_{\oplus}$ planets changes depending on the choice of the faster exponential cutoff. We show the cases where the cutoff rolls after after 1.7 Myr, 1.9 Myr or not at all (i.e. the decay is characterised by only one exponential – the slower one). We see that later cutoffs lead to more divergence.
Figure 6: The curves show the evolution of a pair of $6M_{\oplus}$ planets in the same disk, with the same initial conditions, but with three different values for the time ($t_{c}$) at which the accretion rate starts to rapidly drop off. In the upper left panel, the three black curves represent the inner planet and the three red curves the outer planet. The upper right hand panel shows the evolution of the respective period ratios. In the bottom left, the three panels show the evolution of the resonant angles for the 3:2 resonance (with $t_{c}=1.7$Myr on top, $t_{c}=1.9$ Myr below that and $t_{c}=\infty$ on the bottom). The lower right panels show the corresponding evolution for the 2:1 resonant angles.
@@ 5 Three Planet Systems
Figure 7 shows the simple extension of our model system shown in Figure 2, in which we simply add a third planet, also of mass $3M_{\oplus}$. The initial configuration is chosen so that both pairs are between the 3:2 and 2:1 commensurabilities. The initial evolution is divergent, because the rate of migration is faster closer to the star. The first planet encounters the torque reversal at A, and so the inner pair begins to converge, trapping into the 3:2 resonance at B. This excites the eccentricities of the inner pair to equilibrium values, but the outer planet remains almost circular as it continues to migrate inwards. The resonant interaction slows the inward migration of the middle planet and the outer pair encounters the 3:2 resonance at C. This also briefly disturbs the resonance lock of the inner pair, but both pairs soon settle into resonance again when the middle planet reaches $a_{R}$ (at D). The outer planet continues to migrate inwards, pushing both the inner and middle planets interior to $a_{R}$, until it eventually encounters the torque reversal as well (at E). At this point we have three planets in a resonant chain of 3:2 resonances, moving outwards with the outer planet tracking $a_{R}$. As we have noted above, this will continue until the evolution of $a_{R}$ accelerates and the planets start to diverge (at F). The divergence is not rapid, and the planets eventually freeze out (at G) still in resonance, with a small libration, and moderate separations from commensurability.
Figure 7: Here we show a three planet system, containing three planets, each with mass $3M_{\oplus}$. The evolution of the inner planet is shown in black, the middle planet in red, and the outer planet in blue. In the upper right panel, the red curve refers to period ratio of the inner pair, while the blue curve indicates the period ratio of the outer pair. In the lower right, we show the behaviour of the resonant angles for the 3:2 resonance for both the outer pair (red and blue) and inner pair (red and black). The labels in the upper left indicate important epochs in the evolution of the system, identified as vertical dashed lines in the other panels.
@@
Figure 8 shows the effect of scaling all the masses up by a factor of 2 – i.e. a system containing three planets of mass $6M_{\oplus}$. The inner pair evolves in the same manner as before – converging after the inner planet reaches the torque reversal (A), locking into the 3:2 resonance (B) and finding a new equilibrium when the middle planet reaches the torque reversal (C). However, the stronger gravitational interactions result in the outer pair locking into the 2:1 resonance, at D. Furthermore, the interactions are strong enough to prevent the outer planet from forcing its way inward to $a_{R}$. Thus, the system evolves for a period of time with the middle planet tracking $a_{R}$ and maintaining an inner 3:2 and outer 2:1 resonance.
Figure 8: Here we show a three planet system, containing three planets, each with mass $6M_{\oplus}$. The evolution of the inner planet is shown in black, the middle planet in red, and the outer planet in blue. In the upper right panel, the red curve refers to period ratio of the inner pair, while the blue curve indicates the period ratio of the outer pair. In the lower right, we show the behaviour of the resonant angles for the different first order resonances at play. The upper pair of panels in this quadrant show the evolution of the 3:2 resonance for both the outer pair (red and blue) and inner pair (red and black). The bottom pair of panels show the corresponding evolution for the 2:1 resonance. The labels in the upper left indicate important epochs in the evolution of the system, identified as vertical dashed lines in the other panels.
@@ 5.1 Observational Comparison
These calculations demonstrate that the evolution of systems of multple planets in the protoplanetary disk background calculated by Yu et al. (2023) can produce divergence from resonance and result in final period ratios well away from the commensurabilities expected in a naive application of the resonant locking theory.
In a perfect scenario, we should be able to model the divergence of specific planetary systems with well-defined masses, although the stellar magnetic field value and the evolution of the protoplanetary disk remain quantities that need to be modelled and so can introduce some variations. Of course, the observational reality is less well constrained – most of the known multiplanet systems are detected via the transit method, and most of the planetary masses are small enough to prove challenging to measure with the radial velocity method,
@@
The planetary masses for HIP29442 measured by Damasso et al. (2023) are $4.5\pm 0.3M_{\oplus}$, $5.1\pm 0.4M_{\oplus}$ and $9.6\pm 0.8M_{\oplus}$ for the three planets with orbital periods of 3.538 days (c), 6.429 days (d) and 13.63 days (b). To model this system we choose the scaling of $a_{R}$ so that the innermost planet freezes out at approximately the correct distance. The resulting evolution shows the system enters a 3:2+3:2 resonant chain early and then diverges to final period ratios that are similar to those observed, although slightly smaller. In particular, we note that the outer pair diverged from interior of the 2:1 resonance to outside it.
The masses for the GJ9827 planets (Passegger et al., 2024) are $4.3\pm 0.3M_{\oplus}$, $1.9\pm 0.4M_{\oplus}$ and $3.0\pm 0.6M_{\oplus}$, corresponding to orbital periods of 1.209 days (b), 3.648 days (c) and 6.202 days (d). In this case, we scaled $a_{R}$ so that the middle planet freezes out at the correct distance, and scaled the accretion rate down by a factor of 0.7 to account for the lower stellar mass. With these adjustments our migration model is able to match the outer pair quite well, but the inner planet does not migrate in as far as desired. We have also moved the edge of the magnetospheric cavity out to a value of $a_{R}/2$. This halts the migration of the inward planet once it crosses into the cavity, but the inward migration is still halted by the reversal of the outermost planet when it hits $a_{R}$. Thus, the inner pair mass ratio falls somewhat short of the observed value, but does reproduce the qualitative features of the system in that the inner pair is wider than the 2:1 resonance while the outer pair lies between 3:2 and 2:1.
Of course, most planetary systems do not have masses measured to this accuracy, but our principal goal is not to constrain the histories of individual planets, but to understand the features of the underlying population of planetary systems. Thus, we wish to take a closer look at the results of the Kepler mission, when interpreted in the context of our model.
@@ 5.2 Kepler Analog Systems
As a result, we choose the mass of the middle planet from the results of Chen & Kipping (2018) and then scale the inner and outer planets relative to this using the deterministic component of the mass-radius relation in Chen & Kipping (2017). This allows for some variation in the overall masses, but keeps the internal mass variation consistent with the radius variations. Using these masses we evolve forward realisations of the 107 Kepler planetary systems that contain three confirmed planets, none of which have a radius $>6R_{\oplus}$ (our torque model is not applicable to Jupiter-class planets, which will open gaps in the protoplanetary disk). We wish to focus on the effects of the outspiral and detachment from the disk at late times, so we assume the planets occupy an initially compact configuration, starting the planets at 0.076 AU, 0.107 AU and 0.158 AU, so that the systems lie between the 3:2 and 2:1 resonances.
@@
Figure 10: The solid red points represent the period ratios of inner (horizontal axis) and outer (vertical axis) pairs from our simulations of the set of Kepler triple analoges. The crosses correspond to observed Kepler triples, and the dotted lines indicate the 3:2 and 2:1 commensurabilities. The labelling of the four quadrants A–D are discussed in the text.
We see that our model is capable of producing systems that lie in all four quadrants, as well as a handful that cluster near the resonant chains. This variation is achieved solely as a function of the different masses – all of the red points are generated for the same disk evolution model, initial spatial configurations and $a_{R}$ history. This demonstrates that the mass variations in the Kepler systems are capable of producing a substantial fraction of the overall period ratio variations.
Figure 11 shows how the mass ratios influence the outcome of these models. It shows the distribution of mass ratios for model planetary systems in the four different quadrants. The fact that the systems in Figure 7 and Figure 8 evolve to quadrants A and B respectively, despite starting at the same point (the intersection of the dotted lines in Figure 11), illustrates that total mass plays a role as well, but Figure 11 shows that the mass ratios are an important factor.
@@
Objects in quadrant A tend to have the outermost planet larger than the middle planet, and those that are found far from resonance also tend to have the middle planet larger than the inner one. The planets in quadrant B tend to be found with middle planets larger than inner planets. Those planets in quadrant C tend to exhibit an overall trend of increasing mass outwards, while those in quadrant D tend to have inner pairs with similar mass ratios.
@@
We have already seen examples of evolutionary tracks for quadrants A and B in Figures 7 and 8. Figure 12 shows the example of one that ends up in quadrant C. This is a relatively massive system, with planets of mass 8.1, 9.7 and 10.5 $M_{\oplus}$, a realisation of the masses in the system Kepler-60. It shows that the rapid inward migration causes the inner pair to diverge to the point that they are wider than the 2:1 resonance. When the inner planet halts (at A) this trend reverses and the inner pair lock into the 2:1 resonance (at B). This also slows the inward migration of the middle planet so that the outer pair start to converge as well. However, this happens just before they diverge beyond 2:1, so that the convergence ultimately leads to their trapping in the 3:2 resonance. This transition leads to a a period of dynamical rearrangement, as the inner pair loses its resonant lock at C, before the outer pair actually locks into the 3:2 (at D) and only appears to settle down again once the outer planet reaches the torque reversal (at E).
@@
Figure 12: In this plot, $M_{1}=8.1M_{\oplus}$, $M_{2}=9.7M_{\oplus}$ and $M_{3}=10.4M_{\oplus}$. The colours identify the planets in the same way as in Figure 7. We see that this system starts in the same configuration as the system in Figure 7, but the outer planet diverges so far that it crosses the 2:1 resonance, leaving both pairs wide of the 2:1 commensurability. As shown in the lower right panel, this capture is only temporary but does result in an eccentricity excitation (event G).
@@
After this, the evolution is more stable, with a 2:1+3:2 chain evolving outwards until the system starts to diverge at F. The larger planetary masses means that the system takes longer to decouple, and the outer pair passes through the 2:1 resonance (at G). The system doesn’t trap because the evolution is divergent, but the resonant crossing does produce an eccentricity excitation that doesn’t entirely die away before the system completely freezes out (at H). The end result is that neither pair remains in resonance at the end.
Figure 13: This three planet system is for $M_{1}=4.7M_{\oplus}$, $M_{2}=4.1M_{\oplus}$ and $M_{3}=5.3M_{\oplus}$. The colours refer to the planets in the same order as in Figure 7. In this system, both planets capture into the 3:2 resonance and remain there as each of the planets reach the torque reversal. We see that this system remains in resonance at the end, even though the period ratios have diverged from 3:2 by then.
@@
Figure 13 shows the evolution of a planetary system modelled after Kepler-295, with planetary masses of 4.7, 4.1 and 5.3 $M_{\oplus}$ respectively, representative of a system that ends up in quadrant D. In this case, the evolution is quite similar to that seen in Figure 12, with the inner pair captured into 2:1 after an initial divergence, and the outer pair captured into the 3:2, with a similar brief period of dynamical reshuffling. However, the outer planet is less massive in this case, and the system does not evolve far enough to pass through the 2:1 commensurability, although the outer pair does actually enter the 2:1 resonance by virtue of the large precession at low eccentricities.

### v2

@@ 4.1 Dependance on model parameters
Figure 5 shows the evolution of a pair of $6M_{\oplus}$ planets in three different scenarios. In each case, the accretion disk evolution is the same, but we have scaled the form of $a_{R}$, given in equation (4), by a constant factor. As discussed above, this mimics variations in either the stellar magnetic field strength, or the degree of field compression by the inflow ($n$). Thus the black curve in Figure 5 shows the default case, while the blue curve shows the case where $a_{T}$ is a factor 0.7 smaller (corresponding to a weaker magnetospheric influence) and the red case represents the case where $a_{T}$ is a factor 1.3 larger (corresponding to a stronger magnetospheric influence). In each case, the initial semi-major axes are chosen such that all planets start exterior to $a_{R}$. The blue curves start with semi-major axes 0.063 and 0.08 AU, the black curves with 0.076 and 0.1 AU, and the red curves with 0.084 AU and 0.116 AU. Initial eccentricities are still 0.01 and longitudes are chosen at random.
@@
The final separations also depend on the model for the evolution of the accretion rate and surface density of the disk. Our default model assumes an evolution characterised by two exponentials, chosen to mimic the slow decay of a viscously evolving disk that accelerates once the inner disk has been decoupled from the outer disk by photoevaporation. Figure 6 shows how the evolution of a pair of $6M_{\oplus}$ planets changes depending on the choice of the faster exponential cutoff. Initial conditions for each case are the same as those for Figure 2. We show the evolution where the cutoff rolls over after 1.7 Myr, 1.9 Myr or not at all (i.e. the decay is characterised by only one exponential – the slower one). We see that later cutoffs lead to more divergence.
Figure 6: The curves show the evolution of a pair of $6M_{\oplus}$ planets in the same disk, with the same initial conditions, but with three different values for the time ($t_{c}$) at which the accretion rate starts to rapidly drop off. In the left panel, the three black curves represent the inner planet and the three red curves the outer planet. The right hand panel shows the evolution of the respective period ratios.
@@ 5 Three Planet Systems
Figure 7 shows the simple extension of our model system shown in Figure 2, in which we simply add a third planet, also of mass $3M_{\oplus}$. The initial semi-major axes are now 0.08, 0.107 and 0.149 AU, chosen so that both pairs begin between the 3:2 and 2:1 commensurabilities. Eccentricities and longitudes are chosen as before. The initial evolution is divergent, because the rate of migration is faster closer to the star. The first planet encounters the torque reversal at A, and so the inner pair begins to converge, trapping into the 3:2 resonance at B. This excites the eccentricities of the inner pair to equilibrium values, but the outer planet remains almost circular as it continues to migrate inwards. The resonant interaction slows the inward migration of the middle planet and the outer pair encounters the 3:2 resonance at C. This also briefly disturbs the resonance lock of the inner pair, but both pairs soon settle into resonance again when the middle planet reaches $a_{R}$ (at D), along with a new set of equilibrium eccentricity ratios. The outer planet continues to migrate inwards, pushing both the inner and middle planets interior to $a_{R}$, until it eventually encounters the torque reversal as well (at E). At this point we have three planets in a resonant chain of 3:2 resonances, moving outwards with the outer planet tracking $a_{R}$. As we have noted above, this will continue until the evolution of $a_{R}$ accelerates and the planets start to diverge (at F). The divergence is not rapid, and the planets eventually freeze out (at G) still in resonance, with a small libration, and moderate separations from commensurability.
Figure 7: Here we show a three planet system, containing three planets, each with mass $3M_{\oplus}$. The evolution of the inner planet is shown in black, the middle planet in red, and the outer planet in blue. In the upper right panel, the red curve refers to period ratio of the inner pair, while the blue curve indicates the period ratio of the outer pair. In the lower right, we show the behaviour of the cosines of the resonant angles for the 3:2 resonance for both the outer pair (red and blue) and inner pair (red and black). The labels in the upper left indicate important epochs in the evolution of the system, identified as vertical dashed lines in the other panels.
@@
Figure 8 shows the effect of more massive planets and wider orbits. All the masses are scaled up by a factor of 2 – i.e. a system containing three planets of mass $6M_{\oplus}$. Initial semi-major axes are 0.076, 0.107 and 0.223 AU, so that the outer pair starts wide of the 2:1 resonance. The inner pair evolves in the same manner as before – converging after the inner planet reaches the torque reversal (A), locking into the 3:2 resonance (B) and finding a new equilibrium when the middle planet reaches the torque reversal (C). The outer pair locks into the 2:1 resonance, at D. However, the larger masses mean that the outward torque is stronger and the mutual gravitational interactions are strong enough to prevent the outer planet from forcing its way inward to $a_{R}$. Thus, the system evolves for a period of time with the middle planet tracking $a_{R}$ and maintaining an inner 3:2 and outer 2:1 resonance, with the outer planet being driven outwards despite not having reached $a_{R}$.
Figure 8: Here we show a three planet system, containing three planets, each with mass $6M_{\oplus}$. The evolution of the inner planet is shown in black, the middle planet in red, and the outer planet in blue. In the upper right panel, the red curve refers to period ratio of the inner pair, while the blue curve indicates the period ratio of the outer pair. In the lower right, we show the behaviour of the resonant angles for the different first order resonances at play. The upper panel in this quadrant shows the evolution of the cosine of the resonant angle for the 3:2 resonance for the inner pair (red and black). The bottom panel show the corresponding evolution for the 2:1 resonance between the outer pair of planets. The labels in the upper left indicate important epochs in the evolution of the system, identified as vertical dashed lines in the other panels.
@@ 5.1 Observational Comparison
These calculations demonstrate that the evolution of systems of multiple planets in the protoplanetary disk background calculated by Yu et al. (2023) can produce divergence from resonance and result in final period ratios well away from the commensurabilities expected in a naive application of the resonant locking theory.
In a perfect scenario, we should be able to model the divergence of specific planetary systems with well-defined masses, although the stellar magnetic field value and the evolution of the protoplanetary disk remain quantities that need to be modelled and so can introduce some variations. Of course, the observational reality is less well constrained – most of the known multiplanet systems are detected via the transit method, and most of the planetary masses are small enough to prove challenging to measure with the radial velocity method.
@@
The planetary masses for HIP29442 measured by Damasso et al. (2023) are $4.5\pm 0.3M_{\oplus}$, $5.1\pm 0.4M_{\oplus}$ and $9.6\pm 0.8M_{\oplus}$ for the three planets with orbital periods of 3.538 days (c), 6.429 days (d) and 13.63 days (b). To model this system we choose the scaling of $a_{R}$ so that the innermost planet freezes out at approximately the correct distance. The initial semi-major axes are 0.076, 0.1 and 0.14 AU, with initial eccentricities of 0.01 and random initial longitudes. The resulting evolution shows the system enters a 3:2+3:2 resonant chain early and then diverges to final period ratios that are similar to those observed, although slightly smaller. In particular, we note that the outer pair diverged from interior of the 2:1 resonance to outside it.
The masses for the GJ9827 planets (Passegger et al., 2024) are $4.3\pm 0.3M_{\oplus}$, $1.9\pm 0.4M_{\oplus}$ and $3.0\pm 0.6M_{\oplus}$, corresponding to orbital periods of 1.209 days (b), 3.648 days (c) and 6.202 days (d). In this case, we scaled $a_{R}$ so that the middle planet freezes out at the correct distance, and scaled the accretion rate down by a factor of 0.7 to account for the lower stellar mass. The initial conditions are the same as for HIP29442 except that the inner planet starts with semi-major axis 0.048 AU so that it encounters $a_{R}$ early – since the observed inner planet is very close to the host star. With these adjustments our migration model is able to match the outer pair quite well, but the inner planet does not migrate in as far as desired. We have also moved the edge of the magnetospheric cavity out to a value of $a_{R}/2$. This halts the migration of the inward planet once it crosses into the cavity, but the inward migration is still halted by the reversal of the outermost planet when it hits $a_{R}$. Thus, the inner pair mass ratio falls somewhat short of the observed value, but does reproduce the qualitative features of the system in that the inner pair is wider than the 2:1 resonance while the outer pair lies between 3:2 and 2:1.
Of course, most planetary systems do not have masses measured to this accuracy, but our principal goal is not to constrain the histories of individual planets, but to understand the features of the underlying population of planetary systems. The most representative – in the sense of the most homogenously selected – sample of low mass planets is that derived from the Kepler mission. Thus, we wish to compare our model to the corresponding sample of planetary systems detected with Kepler.
@@ 5.2 Kepler Analog Systems
As a result, we choose the mass of the middle planet from the results of Chen & Kipping (2018) and then scale the inner and outer planets relative to this using the deterministic component of the mass-radius relation in Chen & Kipping (2017). This allows for some variation in the overall masses, but keeps the internal mass variation consistent with the radius variations. Using these masses we evolve forward realisations of the 107 Kepler planetary systems that contain three confirmed planets, none of which have a radius $>6R_{\oplus}$ (our torque model is not applicable to planets of mass $>30M_{\oplus}$ (Yu et al., 2023), which will open gaps in the protoplanetary disk). We wish to focus on the effects of the outspiral and detachment from the disk at late times, so we assume the planets occupy an initially compact configuration, starting the planets at 0.076 AU, 0.107 AU and 0.158 AU, so that the systems lie between the 3:2 and 2:1 resonances. Our model assumes a $1M_{\odot}$ planetary host star, as 85% of the 107 Kepler systems to which we compare orbit FGK stars.
@@
Figure 10: The solid red points represent the period ratios of inner (horizontal axis) and outer (vertical axis) pairs from our simulations of the set of Kepler triple analoges. The red dotted lines indicate the three-body relationship derived from eliminating the longitude of periastron of the middle planet. The crosses correspond to observed Kepler triples, and the dashed lines indicate the 3:2 and 2:1 commensurabilities. The labelling of the four quadrants A–D are discussed in the text.
We see that our model is capable of producing systems that lie in all four quadrants, as well as a handful that cluster near the resonant chains. This variation is achieved solely as a function of the different masses – all of the red points are generated for the same disk evolution model, initial spatial configurations and $a_{R}$ history. This demonstrates that the mass variations in the Kepler systems are capable of producing a substantial fraction of the overall period ratio variations. The red points in each quadrant also show an ordered structure in the immediate vicinity of each of the resonant intersections. These are the result of the fact that a resonant chain of two mean motion resonances also results in a three-body resonance, whose expression is derived by eliminating the longitude of periastron of the middle planet between the two resonances for which it is present (Aksnes, 1988). These are shown as red dotted lines in Figure 10. It should be noted that our equations describe only the first order two-body resonant interactions and so do not contain any zeroth order contributions to the three-body resonance. Nevertheless, the diverging chains preserve the three body relationship well away from the two-body commensurabilities.
@@
Figure 11 shows how the mass ratios influence the outcome of these models. It shows the distribution of mass ratios for model planetary systems in the four different quadrants. Objects in quadrant A tend to have the outermost planet larger than the middle planet, and those that are found far from resonance also tend to have the middle planet larger than the inner one. The planets in quadrant B tend to be found with middle planets larger than inner planets. Those planets in quadrant C tend to exhibit an overall trend of increasing mass outwards, while those in quadrant D tend to have inner pairs with similar mass ratios.
@@
The examples shown in Figure 7 and Figure 8 demonstrate evolutionary pathways that lead to quadrants A and B. Figure 12 shows examples of systems that end up in the other two quadrants, this time drawn from the Kepler analogs. The left hand panel shows a relatively massive system, with planets of mass 8.1, 9.7 and 10.5 $M_{\oplus}$, a realisation of the masses in the system Kepler-60. It shows that the rapid inward migration causes the inner pair to diverge to the point that they are wider than the 2:1 resonance. When the inner planet halts this trend reverses and the inner pair lock into the 2:1 resonance. This also slows the inward migration of the middle planet so that the outer pair start to converge as well. However, this happens just before they diverge beyond 2:1, so that the convergence ultimately leads to their trapping in the 3:2 resonance. This transition leads to a period of dynamical rearrangement, as the inner pair loses its resonant lock, before the outer pair actually locks into the 3:2 resonance and only appears to settle down again once the outer planet reaches the torque reversal.
@@
Figure 12: On the left hand side, $M_{1}=8.1M_{\oplus}$, $M_{2}=9.7M_{\oplus}$ and $M_{3}=10.4M_{\oplus}$. The colors identify the planets in the same way as in Figure 7. We see that this system starts in the same configuration as the system in Figure 7, but the outer planet diverges so far that it crosses the 2:1 resonance, leaving both pairs wide of the 2:1 commensurability and so located in quadrant C of Figure 10. In the lower left, we see that the divergence across the 2:1 resonance results in an eccentricity excitation. On the right hand side, we see the evolution of a three planet system with $M_{1}=4.7M_{\oplus}$, $M_{2}=4.1M_{\oplus}$ and $M_{3}=5.3M_{\oplus}$. In this case the inner pair traps into the 2:1 and the outer pair into the 3:2 resonance. Their eventual divergence leaves the system in quadrant D.
@@
After this, the evolution is more stable, with a 2:1+3:2 chain evolving outwards until eventually the system diverges. The larger planetary masses means that the system takes longer to decouple, and the outer pair passes through the 2:1 resonance. The system doesn’t trap because the evolution is divergent, but the resonant crossing does produce an eccentricity excitation (observed in the lower left panel) that doesn’t entirely die away before the system completely freezes out. The end result is that neither pair remains in resonance at the end.
@@
The right hand side of Figure 12 shows the evolution of a planetary system modelled after Kepler-295, with planetary masses of 4.7, 4.1 and 5.3 $M_{\oplus}$ respectively, representative of a system that ends up in quadrant D. In this case, the evolution is quite similar to that seen in the left hand panel, with the inner pair captured into 2:1 after an initial divergence, and the outer pair captured into the 3:2, with a similar brief period of dynamical reshuffling. However, the outer planet is less massive in this case, and the system does not evolve far enough to pass through the 2:1 commensurability.